# Neuro-Symbolic Control with Large Language Models for Language-Guided Spatial Tasks

### v1

@@ Abstract
Although large language models (LLMs) have recently become effective tools for language-conditioned control in embodied systems, instability, slow convergence, and hallucinated actions continue to limit their direct application to continuous control. A modular neuro-symbolic control framework that clearly distinguishes between low-level motion execution and high-level semantic reasoning is proposed in this work. While a lightweight neural delta controller performs bounded, incremental actions in continuous space, a locally deployed LLM interprets symbolic tasks. We assess the suggested method in a planar manipulation setting with spatial relations between objects specified by language. Numerous tasks and local language models, such as Mistral, Phi, and LLaMA-3.2, are used in extensive experiments to compare LLM-only control, neural-only control, and the suggested LLM+DL framework. In comparison to LLM-only baselines, the results show that the neuro-symbolic integration consistently increases both success rate and efficiency, achieving average step reductions exceeding 70% and speedups of up to $8.83\times$ while remaining robust to language model quality. The suggested framework enhances interpretability, stability, and generalization without any need of reinforcement learning or costly rollouts by controlling the LLM to symbolic outputs and allocating uninterpreted execution to a neural controller trained on artificial geometric data. These outputs show empirically that neuro-symbolic decomposition offers a scalable and principled way to integrate language understanding with ongoing control, this approach promotes the creation of dependable and effective language-guided embodied systems.
@@ 1 Introduction
Recent developments in large language models (LLMs) have shown their exceptional capacity for high-level decision-making, reasoning, and instruction following [2, 1]. These kinds of models are being studied extensively as cognitive engines for robotics and control systems, where they are employed to create the different tasks and plans, choose actions in complex environments, and interpret natural language instructions. However, when used in closed-loop control settings, LLMs show basic limitations despite their expressive power [3]. Particularly, LLM-only control policies typically suffer from hallucinated actions, poor sample efficiency, lack of convergence guarantees, and brittle behavior under feedback, which makes them unreliable for precise spatial manipulation and continuous control tasks [4, 5]. As illustrated in Fig. 1, directly using large language models for continuous control leads to unstable behavior, motivating the proposed neuro-symbolic decomposition.
@@
Nevertheless, directly deploying LLMs as action generators in control loops still remains the problematic [8]. LLM outputs are fundamentally stochastic, sensitive to prompt phrasing, and also not optimized for numerical precision [9]. These type issues become particularly pronounced in spatial control tasks, where the small positional errors can lead to task failure or also towards slow convergence. In term of the result in this scenario, recent studies have highlighted the need for structured interfaces and also the auxiliary control mechanisms to ground LLM reasoning in stable, low-level execution [10].
@@
In the current research work, we propose a neuro-symbolic closed-loop control framework that precisely integrates local large language models with a lightweight type of neural delta controller. So, the LLM is responsible for high-level symbolic reasoning, task interpretation, and also for the goal selection, while the neural controller executes fine-grained corrective actions that ensure fast and stable convergence. By empowering the numerical control to the neural module and restricting the LLM to structured symbolic outputs, the proposed framework clearly mitigates hallucinations, reduces action variance, and significantly improves the sample efficiency.
@@
A fundamental aspect of our proposed approach is the use of locally hosted LLMs, including Mistral, Phi, and LLaMA-3.2. Unlike cloud-based proprietary models, local LLMs facilitate reproducibility, low-latency inference, privacy preservation, and deployment on edge or also resource-constrained systems. Local models, although they are often smaller and less robust than large proprietary counterparts, make them one of the best testbeds for evaluating whether neuro-symbolic integration can compensate for limited model capacity. Our results clearly demonstrate that the proposed hybrid framework consistently improves performance across all evaluated local LLMs, also indicating that the improvements arise from architectural synergy rather than the model scale alone.
@@
We evaluate our method on a suite of spatial reasoning and control tasks in a simulated 2D environment, including relational goals such as right of, left of, above, and below. We compare three settings: LLM-only control, deep learning (DL)-only control, and the proposed LLM+DL hybrid. Extensive experiments show that the hybrid approach achieves substantially higher success rates, reduces the number of control steps by up to an order of magnitude, and yields significant speedups over LLM-only baselines. These improvements are consistent across many different types of LLM architectures and tasks, clearly highlighting the robustness and generality of the proposed framework.
@@ 2.1 Large Language Models for Planning and Control
Large language models have recently been explored as high-level planners and controllers for robotic systems due to their strong reasoning and instruction-following capabilities [11, 12]. Traditional methods demonstrated that LLMs can generate action sequences, code, or symbolic plans from natural-language task descriptions, which enables them to have flexible task specification and zero-shot generalization [13]. Common examples include planning for language-driven frameworks that translate instructions into executable programs or symbolic action graphs.
@@
While promising, LLM-only approaches often suffer from instability, hallucinated actions, and slow convergence due to the lack of grounding in physical dynamics [17]. According to empirical investigations, LLM-only controllers perform poorly on tasks those requiring exact numerical reasoning or spatial accuracy [18], show considerable variance between trials, and are sensitive to prompt phrasing. In most recent work, that has attempted to mitigate these issues through prompt engineering, structured outputs (e.g., JSON schemas), or iterative replanning [19]. However, such methodologies do not essentially address the mismatch among the discrete language reasoning and continuous control execution. But our work explicitly separates symbolic reasoning and numerical control, using the LLM only for high-level task with interpretation while delegating execution to a neural controller trained for stability and precision.
@@ 2.2 Neuro-Symbolic and Hybrid Reasoning Systems
Neuro-symbolic AI aims to combine the comprehensibility and also the compositionality of symbolic reasoning with the scalability and robustness of neural networks [20, 21]. Classical methods especially integrate logic-based systems, rule engines, or symbolic planners with neural perception or control modules [22]. These type of methods have been applied in domains such as visual reasoning, program induction, and task planning [23].
@@ 2.3 Learning-Based Controllers and Delta Control
Prior work has shown that delta-based control significantly improves convergence speed, reduces overshooting, and also enhances robustness in manipulation and navigation tasks [30]. These types of controllers are typically trained using supervised learning or reinforcement learning and also operate efficiently at inference time [31]. Neural network-based trajectory tracking for parallel robots has demonstrated that adaptive learning can also compensate for modeling uncertainties and external disturbances [32].
@@ 2.5 Local and Resource-Constrained Large Language Models
Most of the previous work on LLM-based robotics depends on large proprietary models accessed via cloud APIs [36]. While effective, such approaches raise concerns regarding reproducibility, latency, privacy, and deployment feasibility [37]. Recently, there has been growing interest in local LLMs that can run on consumer-grade hardware, driven by advances in model compression, quantization, and efficient architectures [38].
@@ 3 Methodology
In addition to setting the problem and describing the design of each system component, this part provides the specific range of the suggested neuro-symbolic control that is our proposed framework. The methodology clearly differentiates between symbolic reasoning and continuous control while focusing on the modularity, interpretability, and efficiency.
3.1 Problem Setting and Notation
@@
The objective is to reposition the target marker such that it clearly satisfies a specified spatial relation regarding the reference marker, as described by a natural language instruction.
3.1.1 State Space
@@
where $C\in\mathbb{R}^{+}$ is the maximum coordinate value. The state space is thus defined as the Cartesian product $\mathcal{S}=\mathcal{W}\times\mathcal{W}$.
3.1.2 Task Space
@@
where $\mathbbm{1}[\cdot]$ denotes the indicator function returning 1 if the condition holds and 0 otherwise.
3.1.3 Action Space
@@ 3.2 Overall Neuro-Symbolic Architecture
where the symbolic latent space encoding task semantics denotes by $\mathcal{Z}$.
@@
3.3 Symbolic Reasoning with Local Large Language Models
@@ 3.3.1 LLM Input and Output Constraints
where $\mathcal{Z}=\{0,1,2,3\}$ corresponds to the four canonical spatial relations.
@@ 3.3.2 Multi-Model LLM Integration
without retraining the execution policy. This design enables systematic evaluation of language reasoning quality while keeping the control layer fixed.
@@ 3.4.1 Input Encoding
where $\mathbf{e}_{i}$ denotes the $i$-th standard basis vector, yielding input dimension $d=8$.
3.4.2 Network Architecture
@@
where $\sigma(\cdot)=\max(0,\cdot)$ is the ReLU activation, $\mathbf{W}_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ are weight matrices, and $\mathbf{b}_{l}\in\mathbb{R}^{n_{l}}$ are bias vectors.
@@
3.4.3 Output and State Transition
@@
Delta-based control reduces overshooting and improves stability compared to absolute coordinate prediction. As shown in Fig. 2, the proposed framework decouples symbolic reasoning from continuous control by delegating language understanding to a local LLM and motion execution to a neural delta controller.
@@ 3.5.3 Loss Function
where $\lambda\geq 0$ is the regularization coefficient.
3.5.4 Optimization
@@ 3.6 Closed-Loop Control Strategy
The following process governs the closed-loop operation of the system as follows:
@@ 3.7 Distance-to-Goal Metric
where $\epsilon>0$ ensures numerical stability when $d(\mathbf{s}_{0},\mathcal{T})\approx 0$.
3.8 Baselines and Ablation Configurations
We evaluate three separate system variants to isolate component contributions:
3.8.1 LLM-Only Control
The LLM directly predicts absolute target coordinates:
$(\hat{x}_{r},\hat{y}_{r})=\mathcal{M}_{\text{LLM}}(\mathbf{q}_{t}),$ (30)
also with the displacement computed as follows:
$(\Delta x_{t},\Delta y_{t})=(\hat{x}_{r}-x_{r}^{t},\hat{y}_{r}-y_{r}^{t}).$ (31)
This kind of arrangement is prone to slower convergence, large variation, and hallucinated coordinates.
3.8.2 DL-Only Control
Without using LLM reasoning, the neural controller uses a set, predetermined job encoding:
$(\Delta x_{t},\Delta y_{t})=f_{\theta}(\mathbf{u}_{t}),\quad\text{where }\phi(\mathcal{T})=\text{const}.$ (32)
The neural controller’s ability to maintain control in the absence of symbolic guidance is isolated by this ablation.
3.8.3 LLM+DL (Proposed)
The full neuro-symbolic integration:
$(\Delta x_{t},\Delta y_{t})=f_{\theta}\left(\mathbf{s}_{t},\pi_{\text{sym}}(\mathbf{s}_{t},\mathcal{T})\right),$ (33)
That is combining symbolic task interpretation with neural execution. The proposal technique delivers better robustness across language models, faster convergence, and higher success rates by clearly separating thinking from execution. In addition, this design makes it possible to systematically analyze the quality of LLM reasoning without confusing it with low-level control performance.
@@ 4 Experimental Setup and Results
This section clearly demonstrated the details of the experimental evaluation of the proposed neuro-symbolic control framework. We establish the baseline processes, evaluation metrics, experimental design, and quantitative results. We designed the experiments in such a way that they ensure reproducibility, controlled ablation, and statistically significant comparisons between control strategies and language models.
4.1 Experimental Setup
4.1.1 Environment Configuration
Every experiment is carried out in a planar simulated environment with two identifiable markers: a target marker (red) and a reference marker (blue). The workspace is defined as a bounded square region as follows; $\mathcal{W}=[0,C]^{2}\subset\mathbb{R}^{2}$ with $C=800$ pixels.
At the beginning of each episode, both markers are initialized via independent uniform sampling:
$\mathbf{p}_{r}^{0},\mathbf{p}_{b}^{0}\sim\text{Uniform}(\mathcal{W}).$ (34)
The control objective is to reposition the target marker such that it satisfies a specified spatial relation with respect to the reference marker within a fixed time horizon $T$.

### v2

@@ Abstract
Although large language models (LLMs) have recently become effective tools for language-conditioned control in embodied systems, instability, slow convergence, and hallucinated actions continue to limit their direct application to continuous control. A modular neuro-symbolic control framework that distinguishes between low-level motion execution and high-level semantic reasoning is proposed in this work. While a lightweight neural delta controller performs bounded, incremental actions in continuous space, a locally deployed LLM interprets symbolic tasks. We assess the suggested method in a planar manipulation setting with spatial relations between objects specified by language. Numerous tasks and local language models, such as Mistral, Phi, and LLaMA-3.2, are used in extensive experiments to compare LLM-only control, neural-only control, and the suggested LLM+Deep Learning (LLM+DL) framework. In comparison to LLM-only baselines, the results show that the neuro-symbolic integration consistently increases both success rate and efficiency, achieving average step reductions exceeding 70% and speedups of up to $8.83\times$ while remaining robust to language model quality. The suggested framework enhances interpretability, stability, and generalization without any need of reinforcement learning or costly rollouts by controlling the LLM to symbolic outputs and allocating uninterpreted execution to a neural controller trained on artificial geometric data. These outputs show empirically that neuro-symbolic decomposition offers a scalable and principled way to integrate language understanding with ongoing control, this approach promotes the creation of dependable and effective language-guided embodied systems.
@@ 1 Introduction
Recent developments in large language models (LLMs) have shown their exceptional capacity for high-level decision-making, reasoning, and instruction following [2, 1]. These kinds of models are being studied extensively as cognitive engines for robotics and control systems, where they are employed to create the different tasks and plans, choose actions in complex environments, and interpret natural language instructions. However, when used in closed-loop control settings, LLMs show basic limitations despite their expressive power [3]. Particularly, LLM-only control policies typically suffer from hallucinated actions, poor sample efficiency, lack of convergence guarantees, and brittle behavior under feedback, which makes them unreliable for precise spatial manipulation and continuous control tasks [4, 5]. As illustrated in Figure 1, directly using large language models for continuous control leads to unstable behavior, motivating the proposed neuro-symbolic decomposition.
@@
Nevertheless, directly deploying LLMs as action generators in control loops still remains problematic [8]. LLM outputs are fundamentally stochastic, sensitive to prompt phrasing, and also not optimized for numerical precision [9]. These type issues become particularly pronounced in spatial control tasks, where the small positional errors can lead to task failure or also towards slow convergence. In term of the result in this scenario, recent studies have highlighted the need for structured interfaces and also the auxiliary control mechanisms to ground LLM reasoning in stable, low-level execution [10, 11].
@@
In the current research work, we propose a neuro-symbolic closed-loop control framework that precisely integrates local large language models with a lightweight type of neural delta controller. So, the LLM is responsible for high-level symbolic reasoning, task interpretation, and also for the goal selection, while the neural controller executes fine-grained corrective actions that ensure fast and stable convergence. By empowering the numerical control to the neural module and restricting the LLM to structured symbolic outputs, the proposed framework mitigates hallucinations, reduces action variance, and significantly improves the sample efficiency.
@@
A fundamental aspect of our proposed approach is the use of locally hosted LLMs, including Mistral, Phi, and LLaMA-3.2. Unlike cloud-based proprietary models, local LLMs facilitate reproducibility, low-latency inference, privacy preservation, and deployment on edge or also resource-constrained systems. Local models, although they are often smaller and less robust than large proprietary counterparts, make them one of the best testbeds for evaluating whether neuro-symbolic integration can compensate for limited model capacity. Our results demonstrate that the proposed hybrid framework consistently improves performance across all evaluated local LLMs, also indicating that the improvements arise from architectural synergy rather than the model scale alone.
A fundamental aspect of our proposed approach is the use of locally hosted LLMs, including Mistral, Phi, and LLaMA-3.2. Unlike cloud-based proprietary models, local LLMs facilitate reproducibility, low-latency inference, privacy preservation, and deployment on edge or also resource-constrained systems. Local models, although they are often smaller and less robust than large proprietary counterparts, make them one of the best testbeds for evaluating whether neuro-symbolic integration can compensate for limited model capacity. Our results demonstrate that the proposed hybrid framework consistently improves performance across all evaluated local LLMs, also indicating that the improvements arise from architectural synergy rather than the model scale alone.
@@
We evaluate our method on a suite of spatial reasoning and control tasks in a simulated 2D environment, including relational goals such as right of, left of, above, and below. We compare three settings: LLM-only control, deep learning (DL)-only control, and the proposed LLM+DL hybrid. Extensive experiments show that the hybrid approach achieves substantially higher success rates, reduces the number of control steps by up to an order of magnitude, and yields significant speedups over LLM-only baselines. These improvements are consistent across many different types of LLM architectures and tasks, highlighting the robustness and generality of the proposed framework.
@@ 2.1 Large Language Models for Planning and Control
Large language models have recently been explored as high-level planners and controllers for robotic systems due to their strong reasoning and instruction-following capabilities [12, 13]. Early work demonstrated that LLMs can generate action sequences, code, or symbolic plans from natural-language task descriptions, enabling flexible task specification and zero-shot generalization [14]. Examples include language-driven planning frameworks that translate instructions into executable programs or symbolic action graphs.
@@
While promising, LLM-only approaches often suffer from instability, hallucinated actions, and slow convergence due to the lack of grounding in physical dynamics [18]. According to empirical investigations, LLM-only controllers perform poorly on tasks those requiring exact numerical reasoning or spatial accuracy [19, 20], show considerable variance between trials, and are sensitive to prompt phrasing. In most recent work, that has attempted to mitigate these issues through prompt engineering, structured outputs (e.g., JSON schemas), or iterative replanning [21]. However, such methodologies do not essentially address the mismatch among the discrete language reasoning and continuous control execution. But our work explicitly separates symbolic reasoning and numerical control, using the LLM only for high-level task with interpretation while delegating execution to a neural controller trained for stability and precision.
@@ 2.2 Neuro-Symbolic and Hybrid Reasoning Systems
Neuro-symbolic AI aims to combine the comprehensibility and also the compositionality of symbolic reasoning with the scalability and robustness of neural networks [22, 23]. Classical methods especially integrate logic-based systems, rule engines, or symbolic planners with neural perception or control modules [24]. These type of methods have been applied in domains such as visual reasoning, program induction, and task planning [25, 26].
@@ 2.3 Learning-Based Controllers and Delta Control
Prior work has shown that delta-based control significantly improves convergence speed, reduces overshooting, and also enhances robustness in manipulation and navigation tasks [33]. These types of controllers are typically trained using supervised learning or reinforcement learning and also operate efficiently at inference time [34]. Neural network-based trajectory tracking for parallel robots has demonstrated that adaptive learning can also compensate for modeling uncertainties and external disturbances [35, 36].
@@ 2.5 Local and Resource-Constrained Large Language Models
Most of the previous work on LLM-based robotics depends on large proprietary models accessed via cloud APIs [40, 41]. While effective, such approaches raise concerns regarding reproducibility, latency, privacy, and deployment feasibility [42]. Recently, there has been growing interest in local LLMs that can run on consumer-grade hardware, driven by advances in model compression, quantization, and efficient architectures [43].
@@ 3 Methodology
In addition to setting the problem and describing the design of each system component, this part provides the specific range of the suggested neuro-symbolic control that is our proposed framework. The methodology differentiates between symbolic reasoning and continuous control while focusing on the modularity, interpretability, and efficiency.
3.1 Problem Formulation
@@
The objective is to reposition the target marker such that it satisfies a specified spatial relation regarding the reference marker, as described by a natural language instruction.
@@
where $C\in\mathbb{R}^{+}$ is the maximum coordinate value. The state space is thus defined as the Cartesian product $\mathcal{S}=\mathcal{W}\times\mathcal{W}$. Here, the system state explicitly represents the Cartesian positions of both the target and reference markers within a bounded planar workspace. This formulation provides a minimal yet sufficient representation for spatial relational tasks while avoiding unnecessary state complexity. The bounded workspace constraint ensures that all control actions remain physically valid throughout closed-loop execution.
@@
where $\mathbbm{1}[\cdot]$ denotes the indicator function returning 1 if the condition holds and 0 otherwise. The task space is restricted to a small set of canonical spatial relations, allowing the symbolic reasoning module to operate over a discrete and interpretable output space. The goal predicate $G(\cdot)$ defines task satisfaction using margin-based geometric constraints, which introduces tolerance against small positional errors and supports stable convergence near task boundaries.
@@
    The action space is defined in terms of bounded incremental displacements rather than absolute position commands. This choice prevents large, destabilizing movements and encourages smooth trajectories. The optimization objective prioritizes early task satisfaction, thereby directly minimizing control effort and convergence time in expectation.
@@ 3.2 Overall Neuro-Symbolic Architecture
where the symbolic latent space encoding task semantics denotes by $\mathcal{Z}$. Equation (8) formalizes the neuro-symbolic decomposition central to the proposed framework. High-level symbolic reasoning is handled independently by $\pi_{\mathrm{sym}}$, while continuous execution is delegated to the neural controller $\pi_{\mathrm{neu}}$. This explicit separation allows semantic interpretation and numerical control to be managed by specialized modules, thereby improving control stability, interpretability, and overall robustness.
@@
3.3 Symbolic Reasoning Layer
@@
where $\mathcal{Z}=\{0,1,2,3\}$ corresponds to the four canonical spatial relations. The LLM receives a structured prompt combining the natural language task description with a serialized representation of the current state. Importantly, the LLM output is constrained to a fixed symbolic space through structured parsing. This restriction prevents hallucinated continuous actions and ensures that the language model influences control only through interpretable, discrete decisions.
@@
without retraining the execution policy. This design enables systematic evaluation of language reasoning quality while keeping the control layer fixed. By allowing interchangeable local language models without retraining the neural controller, the framework enables a controlled evaluation of symbolic reasoning quality. This design isolates the effect of the language model from low-level execution dynamics, making performance differences directly attributable to semantic inference rather than control capacity.
@@ 3.4.1 Input Encoding
where $\mathbf{e}_{i}$ denotes the $i$-th standard basis vector, yielding input dimension $d=8$. The neural controller input combines normalized state features with a compact encoding of the symbolic task label. Normalization ensures numerical stability during training and inference, while the task encoding provides explicit semantic conditioning. Both scalar and one-hot encodings are supported, allowing flexibility without altering the controller architecture.
@@
where $\sigma(\cdot)=\max(0,\cdot)$ is the ReLU activation, $\mathbf{W}_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ are weight matrices, and $\mathbf{b}_{l}\in\mathbb{R}^{n_{l}}$ are bias vectors. The neural execution module is implemented as a feedforward network that maps state-task representations to incremental control actions. ReLU activations enable efficient learning of nonlinear control policies, while the fully connected structure ensures low inference latency suitable for closed-loop operation.
@@
The hyperbolic tangent activation enforces strict bounds on the predicted displacements, preventing excessive control actions. This bounded delta formulation reduces overshooting and oscillations commonly observed in LLM-only control, leading to smoother trajectories and more stable convergence.
@@
The state transition updates only the target marker position while enforcing workspace constraints through clipping. This ensures that the system remains within valid spatial limits throughout execution. Combined with bounded actions, this transition model contributes to safe and predictable closed-loop behavior.
Delta-based control reduces overshooting and improves stability compared to absolute coordinate prediction. As shown in Figure 2, the proposed framework decouples symbolic reasoning from continuous control by delegating language understanding to a local LLM and motion execution to a neural delta controller.
@@ 3.5 Training Procedure
where $\lambda\geq 0$ is the regularization coefficient. In the algorithm, the closed-loop interaction between symbolic reasoning and neural execution. At each control step, symbolic guidance is recomputed based on the current state, enabling online correction of reasoning errors. This feedback structure allows the system to recover from imperfect symbolic decisions while maintaining efficient convergence.
@@ 3.6 Closed-Loop Control Strategy
The following process governs the closed-loop operation of the system is shown in the Algorithm 1 below:
@@ 3.7 Distance-to-Goal Metric
where $\epsilon>0$ ensures numerical stability when $d(\mathbf{s}_{0},\mathcal{T})\approx 0$. The distance-to-goal metric provides a continuous measure of task progress that is aligned with the discrete success predicate. Normalization by the initial distance enables fair comparison across episodes with different starting configurations. This metric is used to analyze convergence speed and stability across control strategies.
@@ 4 Experimental Setup and Results
This section demonstrated the details of the experimental evaluation of the proposed neuro-symbolic control framework. We establish the baseline processes, evaluation metrics, experimental design, and quantitative results. We designed the experiments in such a way that they ensure reproducibility, controlled ablation, and statistically significant comparisons between control strategies and language models.
4.1 Environment Configuration
All experiments are conducted in a planar simulated environment containing two point markers: a movable target marker (red) and a static reference marker (blue). The workspace is defined as a bounded square region $\mathcal{W}=[0,C]^{2}\subset\mathbb{R}^{2}$, with $C=800$ pixels.
At the beginning of each episode, the positions of the target and reference markers are independently sampled from a uniform distribution over the workspace. Episodes are executed for a fixed control horizon $T$, and the episode terminates early if the task satisfaction condition is met. All experiments use identical workspace bounds and initialization procedures to ensure consistent evaluation across methods.